OpenStudy (anonymous):

find the vertex -2(x+7)^2+6

OpenStudy (saifoo.khan):

y = 6 at x = -7

OpenStudy (anonymous):

thank you. getting close..!

OpenStudy (saifoo.khan):

:)

OpenStudy (anonymous):

how would I put that as an ordered pair I want to be sure im doing it right

OpenStudy (saifoo.khan):

Alphabetically x comes first, so x first then y.

OpenStudy (anonymous):

ok thank you lol

OpenStudy (saifoo.khan):

uw.

OpenStudy (anonymous):

how do I do this? I want to see the steps b/c im sure i can do it bms x^2+45=12x

OpenStudy (anonymous):

solve for x

OpenStudy (saifoo.khan):

i do them all directly from my Calculator!

OpenStudy (saifoo.khan):

x = 6-3 i x = 6+3 i

OpenStudy (anonymous):

ok nvm lol go a head what is i?

OpenStudy (anonymous):

infinity lol

OpenStudy (saifoo.khan):

its an imaginary number!

OpenStudy (anonymous):

o

OpenStudy (anonymous):

9x^2+6x+1 y and x intercetps

OpenStudy (saifoo.khan):

y int = 1

OpenStudy (anonymous):

1? not 100 ?

OpenStudy (saifoo.khan):

ONE, lol. x int = -1/3

OpenStudy (anonymous):

x^2+4x+4=25 ex and app solutions

OpenStudy (saifoo.khan):

ex? :o

OpenStudy (anonymous):

exact and approximate

OpenStudy (anonymous):

hey here is an easy one checking my work a=3s^2 solve for s wouldnt it be sqrt a/3 or negative a/3

OpenStudy (saifoo.khan):

\[\r =\sqrt{\frac{a}{3}}\]

OpenStudy (anonymous):

ok thats what I put!

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

opening a new box so ic an give another medal

OpenStudy (saifoo.khan):

x = 7, x = 3 y = -21

OpenStudy (saifoo.khan):

Ok. ^^